# Learning multi-modal generative models with permutation-invariant encoders and tighter variational bounds

### v2

@@ Abstract
Devising deep latent variable models for multi-modal data has been a long-standing theme in machine learning research. Multi-modal Variational Autoencoders (VAEs) have been a popular generative model class that learns latent representations that jointly explain multiple modalities. Various objective functions for such models have been suggested, often motivated as lower bounds on the multi-modal data log-likelihood or from information-theoretic considerations. To encode latent variables from different modality subsets, Product-of-Experts (PoE) or Mixture-of-Experts (MoE) aggregation schemes have been routinely used and shown to yield different trade-offs, for instance, regarding their generative quality or consistency across multiple modalities. In this work, we consider a variational bound that can tightly approximate the data log-likelihood. We develop more flexible aggregation schemes that generalize PoE or MoE approaches by combining encoded features from different modalities based on permutation-invariant neural networks. Our numerical experiments illustrate trade-offs for multi-modal variational bounds and various aggregation schemes. We show that tighter variational bounds and more flexible aggregation models can become beneficial when one wants to approximate the true joint distribution over observed modalities and latent variables in identifiable models.
Keywords: Generative Models, Multi-modal Learning, Variational Autoencoders, Permutation-Invariant Neural Networks, Identifiability
@@ 1 Introduction
Multi-modal data sets where each sample has features from distinct sources have grown in recent years. For example, multi-omics data such as genomics, epigenomics, transcriptomics, and metabolomics can provide a more comprehensive understanding of biological systems if multiple modalities are analyzed in an integrative framework (Argelaguet et al., 2018; Lee and van der Schaar, 2021; Minoura et al., 2021). In neuroscience, multi-modal integration of neural activity and behavioral data can help to learn latent neural dynamics (Zhou and Wei, 2020; Schneider et al., 2023). However, annotations or labels in such data sets are often rare, making unsupervised or semi-supervised generative approaches particularly attractive as such methods can be used in these settings to (i) generate data, such as missing modalities, and (ii) learn latent representations that are useful for down-stream analyzes or that are of scientific interest themselves. The availability of heterogeneous data for different modalities promises to learn generalizable representations that can capture shared content across multiple modalities in addition to modality-specific information. A promising class of weakly-supervised generative models is multi-modal VAEs (Suzuki et al., 2016; Wu and Goodman, 2019; Shi et al., 2019; Sutter et al., 2021) that combine information across modalities in an often-shared low-dimensional latent representation. A common route for learning the parameters of latent variable models is via maximization of the marginal data likelihood with various lower bounds thereof, as suggested in previous work.
@@
We consider a set of $M$ random variables $\{X_{1},\ldots,X_{M}\}$ with empirical density $p_{d}$, where each random variable $X_{s}$, $s\in\mathcal{M}=\{1,\dots,M\}$, can be used to model a different data modality taking values in $\mathsf{X}_{s}$. With some abuse of notation, we write $X=\{X_{1},\ldots,X_{M}\}$ and for any subset $\mathcal{S}\subset\mathcal{M}$, we set $X=(X_{\mathcal{S}},X_{\setminus\mathcal{S}})$ for two partitions of the random variables into $X_{\mathcal{S}}=\{X_{s}\}_{s\in\mathcal{S}}$ and $X_{\setminus\mathcal{S}}=\{X_{s}\}_{s\in\mathcal{M}\setminus\mathcal{S}}$. We pursue a latent variable model setup, analogous to uni-modal VAEs (Kingma and Ba, 2014; Rezende et al., 2014). For a latent variable $Z\in\mathsf{Z}$ with prior density $p_{\theta}(z)$, we posit a joint generative model [^1] We usually denote random variables using upper-case letters, and their realizations by the corresponding lower-case letter. We assume throughout that $\mathsf{Z}=\mathbb{R}^{D}$, and that $p_{\theta}(z)$ is a Lebesgue density, although the results can be extended to more general settings such as discrete random variables $Z$ with appropriate adjustments, for instance, regarding the gradient estimators. $p_{\theta}(z,x)=p_{\theta}(z)\prod_{s=1}^{M}p_{\theta}(x_{s}|z)$, where $p_{\theta}(x_{s}|z)$ is commonly referred to as the decoding distribution for modality $s$. Observe that all modalities are independent given the latent variable $z$ shared across all modalities. However, one can introduce modality-specific latent variables by making sparsity assumptions for the decoding distribution.
@@
Popular approaches to train multi-modal models are based on a mixture-based variational bound (Daunhawer et al., 2022; Shi et al., 2019) given by $\mathcal{L}^{\text{Mix}}(\theta,\phi,\beta)=\int\rho(S)\mathcal{L}^{\text{Mix}}_{\mathcal{S}}(x,\theta,\phi,\beta)\mathrm{d}\mathcal{S}$, where
@@
and $\rho$ is some distribution on the power set $\mathcal{P}(\mathcal{M})$ of $\mathcal{M}$ and $\beta>0$. For $\beta=1$, one obtains the bound $\mathcal{L}^{\text{Mix}}_{\mathcal{S}}(x,\theta,\phi,\beta)\leq\log p_{\theta}(x)$. Variations of (1) have been suggested (Sutter et al., 2020), such as by replacing the prior density $p_{\theta}$ in the KL-term by a weighted product of the prior density $p_{\theta}$ and the uni-modal encoding distributions $q_{\phi}(z|x_{s})$, for all $s\in\mathcal{M}$. Maximizing $\mathcal{L}^{\text{Mix}}_{\mathcal{S}}$ can be seen as
minimizing {($\text{minimizing }\left\{\mathcal{H}(X|Z_{\mathcal{S}})+\beta\operatorname{I}_{q_{\phi}}(X_{\mathcal{S}},Z_{\mathcal{S}})=\mathcal{H}(X)-\operatorname{I}_{q_{\phi}}(X,Z_{\mathcal{S}})+\beta\operatorname{I}_{q_{\phi}}(X_{\mathcal{S}},Z_{\mathcal{S}})\right\},$ (2)
@@
where $I_{q}(X,Y)=\int q(x,y)\log\frac{q(x,y)}{q(x)q(y)}$ is the mutual information of random variables $X$ and $Y$ having marginal and joint densities $q$, whilst $\mathcal{H}(X|Y)=-\int q(x,y)\log q(x|y)]\mathrm{d}x\mathrm{d}y$ is the conditional entropy of $X$ given $Y$. We occasionally write $Z_{\mathcal{S}}$ instead of $Z$ to emphasize that $Z$ is conditional on $X_{\mathcal{S}}$ under the encoding density $q_{\phi}$. Likewise, the multi-view variational information bottleneck approach developed in Lee and van der Schaar (2021) for predicting $x_{\setminus\mathcal{S}}$ given $x_{\mathcal{S}}$ can be interpreted as minimizing $-\operatorname{I}_{q_{\phi}}(X_{\setminus\mathcal{S}},Z)+\beta\operatorname{I}_{q_{\phi}}(X_{\mathcal{S}},Z)$. Hwang et al. (2021) suggested a related bound that aims to maximize the reduction of total correlation of $X$ when conditioned on $Z$. Similar bounds have been suggested in Sutter et al. (2020) and Suzuki et al. (2016) by considering different KL-regularisation terms; see also Suzuki and Matsuo (2022). Shi et al. (2020) add a contrastive term to the maximum likelihood objective and minimize $-\log p_{\theta}(x)-\beta\operatorname{I}_{p_{\theta}}(X_{\mathcal{S}},X_{\setminus\mathcal{S}})$.
@@
  • Product of Experts (PoE), see Wu and Goodman (2018),
@@
This paper contributes (i) a new variational bound as an approximate lower bound on the multi-modal log-likelihood (LLH). We avoid a limitation of mixture-based bounds (1), which may not provide tight lower bounds on the joint LLH if there is considerable modality-specific variation (Daunhawer et al., 2022), even for flexible encoding distributions. The novel variational bound contains a lower bound of the marginal LLH $\log p_{\theta}(x_{\mathcal{S}})$ and a term approximating the conditional $\log p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$ for any choice of $\mathcal{S}\in\mathcal{P}(\mathcal{M})$, provided that we can learn a flexible multi-modal encoding distribution. This paper then contributes (ii) new multi-modal aggregation schemes that yield more expressive multi-modal encoding distributions when compared to MoEs or PoEs. These schemes are motivated by the flexibility of permutation-invariant (PI) architectures such as DeepSets (Zaheer et al., 2017) or attention models (Vaswani et al., 2017; Lee et al., 2019). We illustrate that these innovations (iii) are beneficial when learning identifiable models, aided by using flexible prior and encoding distributions consisting of mixtures and (iv) yield higher LLH in experiments.
@@
Canonical Correlation Analysis (Hotelling, 1936; Bach and Jordan, 2005) is a classical approach for multi-modal data that aims to find projections of two modalities by maximally correlating and has been extended to include more than two modalities (Archambeau and Bach, 2008; Tenenhaus and Tenenhaus, 2011) or to allow for non-linear transformations (Akaho, 2001; Hardoon et al., 2004; Wang et al., 2015; Karami and Schuurmans, 2021). Probabilistic CCA can also be seen as multi-battery factor analysis (MBFA) (Browne, 1980; Klami et al., 2013), wherein a shared latent variable models the variation common to all modalities with modality-specific latent variables capturing the remaining variation. Likewise, latent factor regression or classification models (Stock and Watson, 2002) assume that observed features and response are driven jointly by a latent variable. Vedantam et al. (2018) considered a tiple-ELBO for two modalities, while Sutter et al. (2021) introduced a generalized variational bound that involves a summation over all modality subsets. A series of work has developed multi-modal VAEs based on shared and private latent variables (Wang et al., 2016; Lee and Pavlovic, 2021; Lyu and Fu, 2022; Lyu et al., 2021; Vasco et al., 2022; Palumbo et al., 2023). Tsai et al. (2019) proposed a hybrid generative-discriminative objective and minimized an approximation of the Wasserstein distance between the generated and observed multi-modal data. Joy et al. (2021) consider a semi-supervised setup of two modalities that requires no explicit multi-modal aggregation function. Extending the Info-Max principle (Linsker, 1988), maximizing mutual information $\operatorname{I}_{q}(g_{1}(X_{1}),g(X_{2}))\leq\operatorname{I}_{q}((X_{1},X_{2}),(Z_{1},Z_{2}))$ based on representations $Z_{s}=g_{s}(X_{s})$ for modality-specific encoders $g_{s}$ from two modalities has been a motivation for approaches based on (symmetrized) contrastive objectives (Tian et al., 2020; Zhang et al., 2022c; Daunhawer et al., 2023) such as InfoNCE (Oord et al., 2018; Poole et al., 2019; Wang and Isola, 2020) as a variational lower bound on the mutual information between $Z_{1}$ and $Z_{2}$. Recent work (Bounoua et al., 2023; Bao et al., 2023) considered score-based diffusion models on auto-encoded private latent variables.
2 A tighter variational bound with arbitrary modality masking

### v3

# Learning multi-modal generative models with permutation-invariant encoders and tighter variational objectives

@@ Abstract
Devising deep latent variable models for multi-modal data has been a long-standing theme in machine learning research. Multi-modal Variational Autoencoders (VAEs) have been a popular generative model class that learns latent representations that jointly explain multiple modalities. Various objective functions for such models have been suggested, often motivated as lower bounds on the multi-modal data log-likelihood or from information-theoretic considerations. To encode latent variables from different modality subsets, Product-of-Experts (PoE) or Mixture-of-Experts (MoE) aggregation schemes have been routinely used and shown to yield different trade-offs, for instance, regarding their generative quality or consistency across multiple modalities. In this work, we consider a variational objective that can tightly approximate the data log-likelihood. We develop more flexible aggregation schemes that avoid the inductive biases in PoE or MoE approaches by combining encoded features from different modalities based on permutation-invariant neural networks. Our numerical experiments illustrate trade-offs for multi-modal variational objectives and various aggregation schemes. We show that our variational objective and more flexible aggregation models can become beneficial when one wants to approximate the true joint distribution over observed modalities and latent variables in identifiable models.
@@ 1 Introduction
Multi-modal data sets where each sample has features from distinct sources have grown in recent years. For example, multi-omics data such as genomics, epigenomics, transcriptomics, and metabolomics can provide a more comprehensive understanding of biological systems if multiple modalities are analyzed in an integrative framework [Argelaguet et al., 2018, Lee and van der Schaar, 2021, Minoura et al., 2021]. In neuroscience, multi-modal integration of neural activity and behavioral data can help to learn latent neural dynamics [Zhou and Wei, 2020, Schneider et al., 2023]. However, annotations or labels in such data sets are often rare, making unsupervised or semi-supervised generative approaches particularly attractive as such methods can be used in these settings to (i) generate data, such as missing modalities, and (ii) learn latent representations that are useful for down-stream analyzes or that are of scientific interest themselves. The availability of heterogeneous data for different modalities promises to learn generalizable representations that can capture shared content across multiple modalities in addition to modality-specific information. A promising class of weakly-supervised generative models is multi-modal VAEs [Suzuki et al., 2016, Wu and Goodman, 2019, Shi et al., 2019, Sutter et al., 2021] that combine information across modalities in an often-shared low-dimensional latent representation. A common route for learning the parameters of latent variable models is via maximization of the marginal data likelihood with various lower bounds thereof, as suggested in previous work.
@@
We consider a set of $M$ random variables $\{X_{1},\ldots,X_{M}\}$ with empirical density $p_{d}$, where each random variable $X_{s}$, $s\in\mathcal{M}=\{1,\dots,M\}$, can be used to model a different data modality taking values in $\mathsf{X}_{s}$. With some abuse of notation, we write $X=\{X_{1},\ldots,X_{M}\}$ and for any subset $\mathcal{S}\subset\mathcal{M}$, we set $X=(X_{\mathcal{S}},X_{\setminus\mathcal{S}})$ for two partitions of the random variables into $X_{\mathcal{S}}=\{X_{s}\}_{s\in\mathcal{S}}$ and $X_{\setminus\mathcal{S}}=\{X_{s}\}_{s\in\mathcal{M}\setminus\mathcal{S}}$. We pursue a latent variable model setup, analogous to uni-modal VAEs [Kingma and Ba, 2014, Rezende et al., 2014]. For a latent variable $Z\in\mathsf{Z}$ with prior density $p_{\theta}(z)$, we posit a joint generative model [^1] We usually denote random variables using upper-case letters, and their realizations by the corresponding lower-case letter. We assume throughout that $\mathsf{Z}=\mathbb{R}^{D}$, and that $p_{\theta}(z)$ is a Lebesgue density, although the results can be extended to more general settings such as discrete random variables $Z$ with appropriate adjustments, for instance, regarding the gradient estimators. $p_{\theta}(z,x)=p_{\theta}(z)\prod_{s=1}^{M}p_{\theta}(x_{s}|z)$, where $p_{\theta}(x_{s}|z)$ is commonly referred to as the decoding distribution for modality $s$. Observe that all modalities are independent given the latent variable $z$ shared across all modalities. However, one can introduce modality-specific latent variables by making sparsity assumptions for the decoding distribution. Intuitively, this conditional independence assumption means that the latent variable $Z$ captures all unobserved factors shared by the modalities.
Figure 1: Reconstruction or cross-prediction of modalities in 1(a) and 1(b) for a mixture-based bound and our objective, respectively. The mixture-based bound resorts to a single latent variable $Z\sim q_{\phi}(\cdot|x_{\mathcal{S}})$ that encodes information from a modality subset $x_{\mathcal{S}}$ and is trained to reconstruct the conditioning modalities $x_{\mathcal{S}}$, as well as to predict the masked modalities $x_{\setminus\mathcal{S}}$. Our objective relies on two latent variables $Z_{\mathcal{S}}\sim q_{\phi}(\cdot|x_{\mathcal{S}})$ and $Z_{\mathcal{M}}\sim q_{\phi}(\cdot|x_{\mathcal{S}},x_{\setminus\mathcal{S}})$, where $Z_{\mathcal{S}}$ is learned to reconstruct all its conditioning modalities, with $Z_{\mathcal{M}}$ learned to reconstruct the remaining modalities. KL regularization terms in 1(c) and 1(d) for a mixture-based bound and our objective, respectively. The mixture-based bound aims to minimize the KL divergence between the encoding distribution given a modality subset $x_{\mathcal{S}}$ and a prior distribution. Our objective additionally aims to minimize the KL divergence between the encoding distribution given all modalities relative to the encoding distribution of a modality subset $x_{\mathcal{S}}$.
@@
Popular approaches to train multi-modal models are based on a mixture-based variational bound [Daunhawer et al., 2022, Shi et al., 2019] given by $\mathcal{L}^{\text{Mix}}(\theta,\phi,\beta)=\int\rho(S)\mathcal{L}^{\text{Mix}}_{\mathcal{S}}(x,\theta,\phi,\beta)\mathrm{d}\mathcal{S}$, where
@@
and $\rho$ is some distribution on the power set $\mathcal{P}(\mathcal{M})$ of $\mathcal{M}$ and $\beta>0$. For $\beta=1$, one obtains the bound $\mathcal{L}^{\text{Mix}}_{\mathcal{S}}(x,\theta,\phi,\beta)\leq\log p_{\theta}(x)$. However, as shown in Daunhawer et al. [2022], there is a gap between the variational bound and the log-likelihood given by the conditional entropies that cannot be reduced even for flexible encoding distributions. More precisely, it holds that
$\int p_{d}(x)\log p_{\theta}(x)\mathrm{d}x\geq\int p_{d}(x)\mathcal{L}^{\text{Mix}}(x,\theta,\phi,1)\mathrm{d}x+\int\rho(\mathcal{S})\mathcal{H}(p_{d}(X_{\setminus\mathcal{S}}|X_{\mathcal{S}}))\mathrm{d}\mathcal{S},$
where $\mathcal{H}(p_{d}(X_{\setminus\mathcal{S}}|X_{\mathcal{S}}))$ is the entropy of the conditional data distributions. Intuitively, in (1), one tries to reconstruct or predict all modalities from incomplete information using only the modalities $\mathcal{S}$, which leads to learning an inexact, average prediction [Daunhawer et al., 2022]. In particular, it cannot reliably predict modality-specific information that is not shared with other modality subsets, as measured by the conditional entropies $\mathcal{H}(p_{d}(X_{\setminus\mathcal{S}}|X_{\mathcal{S}}))$. For an illustration, see Figure 1(a), where the latent variable $Z$ encodes information from a text and audio modality and is tasked to both reconstruct the text and audio modalities and to predict an unobserved image. Information that is specific to the image modality only thus cannot be recovered. A related observation has been made for Masked AutoEncoders [He et al., 2022] that correspond to the limiting case $\beta\to 0$, where the cross-reconstruction objective leads to learning features that are invariant to the masking of modalities [Kong and Zhang, 2023] and allows for the recovery of latent variables that represent maximally shared information between the unmasked and masked modality [Kong et al., 2023].
We will illustrate that maximizing $\mathcal{L}^{\text{Mix}}_{\mathcal{S}}$ can be interpreted as the information-theoretic objective of
maximizing $\text{maximizing }\left\{\hat{\operatorname{I}}^{\text{lb}}_{q_{\phi}}(X,Z_{\mathcal{S}})-\beta\hat{\operatorname{I}}^{\text{ub}}_{q_{\phi}}(X_{\mathcal{S}},Z_{\mathcal{S}})\right\},$ (2)
@@
where $\hat{\operatorname{I}}^{\text{ub}}_{q}$ and $\hat{\operatorname{I}}^{\text{lb}}_{q}$ are variational upper, respectively, lower bounds of the corresponding mutual information $I_{q}(X,Y)=\int q(x,y)\log\frac{q(x,y)}{q(x)q(y)}\mathrm{d}x\mathrm{d}y$ of random variables $X$ and $Y$ having marginal and joint densities $q$. In order to emphasize that the latent variable $Z$ is conditional on $X_{\mathcal{S}}$ under the encoding density $q_{\phi}$, we write $Z_{\mathcal{S}}$ instead of $Z$. Variations of (1) have been suggested [Sutter et al., 2020], such as by replacing the prior density $p_{\theta}$ in the KL-term by a weighted product of the prior density $p_{\theta}$ and the uni-modal encoding distributions $q_{\phi}(z|x_{s})$, for all $s\in\mathcal{M}$. Likewise, the multi-view variational information bottleneck approach developed in Lee and van der Schaar [2021] for predicting $X_{\setminus\mathcal{S}}$ given $X_{\mathcal{S}}$ can be interpreted as maximizing $\hat{\operatorname{I}}^{\text{lb}}_{q_{\phi}}(X_{\setminus\mathcal{S}},Z_{\mathcal{S}})-\beta\hat{\operatorname{I}}^{\text{ub}}_{q_{\phi}}(X_{\mathcal{S}},Z_{\mathcal{S}})$. Hwang et al. [2021] suggested a related bound that aims to maximize the reduction of total correlation of $X$ when conditioned on a latent variable. Similar bounds have been suggested in Sutter et al. [2020] and Suzuki et al. [2016] by considering different KL-regularisation terms; see also Suzuki and Matsuo [2022]. Shi et al. [2020] add a contrastive estimate $\hat{\operatorname{I}}_{p_{\theta}}$ of the point-wise mutual information to the maximum likelihood objective and minimize $-\log p_{\theta}(x)-\beta\hat{\operatorname{I}}_{p_{\theta}}(x_{\mathcal{S}},x_{\setminus\mathcal{S}})$. Optimizing variational bounds of different mutual information terms such as (2) yield latent representations that have different trade-offs in terms of either (i) reconstruction or (ii) cross-prediction of multi-modal data from a rate-distortion viewpoint [Alemi et al., 2018].
@@
  • Product of Experts (PoE), see Wu and Goodman [2018],
@@
Figure 2(a)-2(b) illustrates these previously considered aggregation schemes. While these schemes do not require learning the aggregation function, we introduce aggregation schemes that, as illustrated in 2(c)-2(d), involve learning additional neural network parameters that specify a learnable aggregation function.
Figure 2: Illustration of multi-modal aggregation schemes. All encoding schemes first apply modality-specific encoders to each individual modality. A PoE model (2(a)) aggregates the outputs from the modality-specific encoders into a single Gaussian distribution that results from a multiplication of the corresponding uni-modal Gaussian densities. An MoE model (2(b)) assumes an equally weighted Gaussian mixture distribution comprised of the uni-modal Gaussian densities. Our new aggregation schemes allow for learning permutation-invariant fusion models: A Sum-Pooling or Deep Set model (2(c)) applies the same function $g$ to the encoded features $h_{s}$, $s\in\{T,A,I\}$, before summing them up and using a non-linear projection $\rho$ to the parameters of a Gaussian distribution. A Self-Attention model (2(d)) differs from the Sum-Pooling approach by applying self-attention layers or transformer layers before summing up the features, thereby accounting for pairwise interactions between the encoded modalities. Our newly introduced schemes allow for encoding only a modality subset by using standard masking operations.
@@
This paper contributes (i) a new variational objective as an approximation of a lower bound on the multi-modal log-likelihood (LLH). We avoid a limitation of mixture-based bounds (1), which may not provide tight lower bounds on the joint LLH if there is considerable modality-specific variation [Daunhawer et al., 2022], even for flexible encoding distributions. The novel variational objective contains a lower bound of the marginal LLH $\log p_{\theta}(x_{\mathcal{S}})$ and a term approximating the conditional $\log p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$ for any choice of $\mathcal{S}\in\mathcal{P}(\mathcal{M})$, provided that we can learn a flexible multi-modal encoding distribution. This paper then contributes (ii) new multi-modal aggregation schemes that yield more expressive multi-modal encoding distributions when compared to MoEs or PoEs. These schemes are motivated by the flexibility of permutation-invariant (PI) architectures such as DeepSets [Zaheer et al., 2017] or attention models [Vaswani et al., 2017, Lee et al., 2019]. We illustrate that these innovations (iii) are beneficial when learning identifiable models, aided by using flexible prior and encoding distributions consisting of mixtures, and (iv) yield higher LLH in experiments.
@@
Canonical Correlation Analysis [Hotelling, 1936, Bach and Jordan, 2005] is a classical approach for multi-modal data that aims to find projections of two modalities by maximally correlating and has been extended to include more than two modalities [Archambeau and Bach, 2008, Tenenhaus and Tenenhaus, 2011] or to allow for non-linear transformations [Akaho, 2001, Hardoon et al., 2004, Wang et al., 2015, Karami and Schuurmans, 2021]. Probabilistic CCA can also be seen as multi-battery factor analysis (MBFA) [Browne, 1980, Klami et al., 2013], wherein a shared latent variable models the variation common to all modalities with modality-specific latent variables capturing the remaining variation. Likewise, latent factor regression or classification models [Stock and Watson, 2002] assume that observed features and response are driven jointly by a latent variable. Vedantam et al. [2018] considered a tiple-ELBO for two modalities, while Sutter et al. [2021] introduced a generalized variational bound that involves a summation over all modality subsets. A series of work has developed multi-modal VAEs based on shared and private latent variables [Wang et al., 2016, Lee and Pavlovic, 2021, Lyu and Fu, 2022, Lyu et al., 2021, Vasco et al., 2022, Palumbo et al., 2023]. Tsai et al. [2019a] proposed a hybrid generative-discriminative objective and minimized an approximation of the Wasserstein distance between the generated and observed multi-modal data. Joy et al. [2021] consider a semi-supervised setup of two modalities that requires no explicit multi-modal aggregation function. Extending the Info-Max principle [Linsker, 1988], maximizing mutual information $\operatorname{I}_{q}(g_{1}(X_{1}),g(X_{2}))\leq\operatorname{I}_{q}((X_{1},X_{2}),(Z_{1},Z_{2}))$ based on representations $Z_{s}=g_{s}(X_{s})$ for modality-specific encoders $g_{s}$ from two modalities has been a motivation for approaches based on (symmetrized) contrastive objectives [Tian et al., 2020, Zhang et al., 2022c, Daunhawer et al., 2023] such as InfoNCE [Oord et al., 2018, Poole et al., 2019, Wang and Isola, 2020] as a variational lower bound on the mutual information between $Z_{1}$ and $Z_{2}$. Recent work [Bounoua et al., 2023, Bao et al., 2023] considered score-based diffusion models on auto-encoded private latent variables.
2 A tighter variational objective with arbitrary modality masking